# Matijevic-Roberts type theorems, Rees rings and associated graded rings

### v1

@@ 1. Introduction
Historically, Rees rings and associated graded rings of Noetherian rings were studied in connection with the problem of resolving singularities of Noetherian schemes by blowing up closed subschemes. The aim of the present article is to study these objects from the ring-theoretic viewpoint. One often encounters a situation where many notable ring-theoretic properties can be derived by studying associated graded rings. Geometrically, associated graded rings correspond to exceptional divisors of a blown-up scheme. Let $A$ be a Noetherian ring with an ideal $I\subseteq A$ and let $\mathscr{R}_{+}(I)$ be its Rees ring. For example, Barshay proved that if $A$ is Cohen-Macaulay and $I$ is generated by a regular sequence, then $\mathscr{R}_{+}(I)$ is also Cohen-Macaulay in [3]. Let $G(I)$ be the associated graded ring of $A$ with respect to $I$. Then Goto and Shimoda characterized the Cohen-Macaulayness/Gorensteinness of $\mathscr{R}_{+}(I)$ in terms of the corresponding property of $G(I)$ and its attached $a$-invariant in [9]. Conversely, one can ask how the singularities of $G(I)$ affect the singularities of $A$. Suppose that $\mathcal{P}$ is one of the following properties: Cohen-Macaulay, Gorenstein, reduced, domain, or normal. Then under a very mind condition, if $G(I)$ has $\mathcal{P}$, then so does $A$ (see [6, Theorem 4.5.7, Theorem 4.5.8 and Theorem 4.5.9] for the proofs). We use Rees rings, extended Rees rings, and associated rings to study the relationship among three notable problems in commutative algebra, which are Lifting Problem, Ascent/descent Problem in associated graded rings, and Matijevic-Roberts type Problem. Our guiding principle is explained by Proposition 3.1 and Proposition 3.5. As a test example, we consider the case $\mathcal{P}=\rm{weakly~{}normal,~{}seminormal}$. See Theorem 3.6 and Theorem 3.8.
@@ 2. Notation
As we already mentioned in the introduction, a remarkable aspect of associated graded rings is that one can often derive a certain property of $A$ from $G(I)$. For instance, it is known that if the associated graded ring with respect to some proper ideal of $A$ is Cohen-Macaulay, then $A$ is also Cohen-Macaulay. A similar statement holds for the Gorenstein property (see [14, Proposition 1.2]). Let us recall some definitions. See a survey paper [20] for other equivalent definitions of seminormality and weak normality.
@@
Let $A$ be a commutative Noetherian ring. Consider the following conditions.
@@
A ring $A$ which satisfies the condition $\rm{(i)}$ is called seminormal, and which satisfies both conditions $\rm{(i)}$ and $\rm{(ii)}$ is called weakly normal.
Let us recall the notion of algebra retracts, following Costa.
Proposition 2.2 ([7, Proposition 1.1]).
Let $f:R\to S$ be an extension of commutative rings. Then the following conditions are equivalent.
  1. (1)
    There is a ring map $g:S\to S$ such that $g\circ g=g$ and $g(S)=R$.
  2. (2)
    There is a ring map $h:S\to R$ such that $h\circ f=\operatorname{id}_{R}$.
  3. (3)
    There is a decomposition of rings $S\cong R\oplus I$ for some ideal $I\subseteq S$ and this decomposition is $R$-linear.
Definition 2.3.
Let $f:R\to S$ be a ring extension and let the notation be as in Proposition 2.2. Then we say that $h$ is a retraction of $S$ and $R$ is an algebra retract (or simply retract) of $S$.
We should notice that the property of $R$ being an algebra retract of $S$ is stronger than $R$ being an $R$-direct summand of $S$ as an $R$-module. Indeed, the condition of an algebra retract gives a control on a certain ring-theoretic property of $R$ inherited from $S$. For example, if $S$ is a unique factorization domain, then so is $R$ (see [7, Proposition 1.8]).
@@ 1. $\bullet$
    Let $A\in\mathbf{D}$. Then any finitely generated $A$-algebra belongs $\mathbf{D}$, the localization of $A$ with respect to any multiplicative subset of $A$ belongs to $\mathbf{D}$, and any subring $B\subseteq A$ such that $B$ is an $A$-algebra retract belongs to $\mathbf{D}$. Let $I\subseteq A$ be an ideal. Then the (extended) Rees ring of $A$ with respect to $I$ belongs to $\mathbf{D}$.
We note that if $A$ is an algebra retract of a Noetherian ring, then $A$ is also Noetherian by Proposition 2.2. Henceforth, we fix $\mathbf{D}$ as above and consider a ring-theoretic property $\mathcal{P}$ which is defined on any object of $\mathbf{D}$. We consider the following properties on the objects from $\mathbf{D}$.
@@
    If $A$ has $\mathcal{P}$ and $B$ is an algebra retract of $A$, then $B$ has $\mathcal{P}$.
@@
by letting $X\mapsto\overline{y}\in yA/y^{2}A$. Then $\phi$ is an isomorphism in view of [12, Theorem 16.2]. Since $A/yA$ has $\mathcal{P}$ by assumption, the condition $(\bf{P2})$ allows us to say that $(A/yA)[X]$ has $\mathcal{P}$. By the graded isomorphism $\phi$, we see that the associated graded ring $G(yA)$ has $\mathcal{P}$. By virtue of assumption $(\bf{Gr})$, it follows that $A$ has $\mathcal{P}$. In other words, $(\bf{Lift})$ is solved. ∎
We recall the following result (see [15, Proposition I.7.4]).
@@
An application of Proposition 3.1 is a new proof of Theorem 3.2, which was already mentioned in [8, Remark 3.8]. However, it seems that most existing research articles study $(\bf{Lift})$ extensively, while $(\bf{Gr})$ is less tractable. As a side topic, we prove an ideal-theoretic analogue of Theorem 3.2 under some conditions.
@@
which is the Rees ring $(A/yA)[(I/(I\cap yA))t]\subseteq(A/yA)[t]$. By assumption, $I^{n}/(I^{n}\cap yA)\subseteq A/yA$ is a normal ideal and $A/yA$ is integrally closed, so $(A/yA)[(I/(I\cap yA))t]$ is integrally closed in a normal domain $(A/yA)[t]$ in view of [18, Proposition 5.2.1]. Hence $\mathscr{R}_{+}(I)/y\mathscr{R}_{+}(I)$ is an integrally closed domain. Let $\mathfrak{M}:=\mathfrak{m}\oplus\bigoplus_{n\geq 1}I^{n}t^{n}$, which is the unique graded maximal ideal of $\mathscr{R}_{+}(I)$. Since $y\in\mathfrak{M}$, it follows that $\mathscr{R}_{+}(I)_{\mathfrak{M}}$ is a local normal domain. By applying [16, Proposition 2.2], we find that $\mathscr{R}_{+}(I)$ is an integrally closed domain. In particular, this ring is integrally closed in $A[t]$, which gives that $I$ is a normal ideal by [18, Proposition 5.2.1]. ∎
@@
where $\mathscr{R}(I)=A[It,t^{-1}]$ is the extended Rees ring. Note that $\mathscr{R}(I)$ has a unique graded maximal ideal $\mathfrak{M}=\mathfrak{m}\mathscr{R}(I)+It\mathscr{R}(I)+t^{-1}\mathscr{R}(I)$ containing $t^{-1}$. Moreover, $t^{-1}$ is a nonzero divisor of $\mathscr{R}(I)$. By $(\bf{Lift})$ and $(\bf{P1})$, we find that the localization $\mathscr{R}(I)_{\mathfrak{M}}$ has $\mathcal{P}$. Then $(\bf{MR})$ shows that $\mathscr{R}(I)$ has $\mathcal{P}$. Since $A$ is a degree-zero part of $\mathscr{R}(I)$, thus an algebra retract. Hence the condition $(\bf{P3})$ allows us to conclude that $A$ has $\mathcal{P}$, as desired. ∎
@@
$(1)$: Since seminormality is a local property in view of [17, Proposition 3.7], the “only if” part is clear. It suffices to prove the “if” part. So assume that $A_{\mathfrak{m}}$ is seminormal for every graded maximal ideal $\mathfrak{m}$. Again by [17, Proposition 3.7], it suffices to show that $A_{\mathfrak{p}}$ is seminormal for every maximal ideal $\mathfrak{p}\subseteq A$. Let $\mathfrak{p}^{*}$ be the ideal generated by all homogeneous elements contained in $\mathfrak{p}$. Then it follows from [6, Lemma 1.5.6] that $\mathfrak{p}^{*}$ is a prime ideal. Let $A_{(\mathfrak{p})}$ and $A_{(\mathfrak{p}^{*})}$ be the homogeneous localizations as in [6, p.31]. Then by the construction, we have $A_{(\mathfrak{p}^{*})}=A_{(\mathfrak{p})}$. Since $\mathfrak{p}^{*}$ is graded, there is a graded maximal ideal $\mathfrak{m}$ such that $\mathfrak{p}^{*}\subseteq\mathfrak{m}$. Then we have a localization map $A_{(\mathfrak{m})}\to A_{(\mathfrak{p}^{*})}=A_{(\mathfrak{p})}$, which extends to
@@
However, as the only graded maximal ideal of $A$ is $\mathfrak{m}$ and $b$ is a homogeneous element not contained in $\mathfrak{m}$, we find that $b$ is a unit element of $A$. That is, we get $x\in A$. By applying [11, Proposition 2.4] (see also [1, Theorem 2]), $A$ is seminormal, as desired.
$(2)$: The weak normality is a local property by [21, Corollary 2]. So as in the first case, we may assume that $R$ is a $\mathbb{Z}$-graded ring with a unique maximal ideal $\mathfrak{m}$ and $A_{\mathfrak{m}}$ is weakly normal. Keep the notation as in $(1)$. In view of [19, Proposition 3.1], it suffices to show that if a homogeneous element $x\in\overline{A}$ satisfies $x^{p}\in A$ and $px\in A$ for some prime $p$, then $x\in A$. By the weak normality of $A_{\mathfrak{m}}$, it follows that $\frac{x}{1}\in A_{\mathfrak{m}}$ and one can proceed as in the first case to finish the proof. ∎
@@
While the seminormal case due to Heitmann is found in [10], the weakly normal case due to Murayama is found in [13, Proposition 4.10], where a new proof of the seminormal case is also given. Recently, even another innovative approach is discovered in [4]. The following result gives a realization of Proposition 3.5. The corresponding statement in the normal case has been already known. See for instance [6, Theorem 4.5.9].
@@
$(1)$: First, we prove that the extended Rees ring $\mathscr{R}(I)$ is seminormal. Recall that $G(I)\cong\mathscr{R}(I)/t^{-1}\mathscr{R}(I)$ and $t^{-1}$ is a nonzero divisor of $\mathscr{R}(I)$. Since $A$ is local, $\mathscr{R}(I)$ has a unique graded maximal ideal $\mathfrak{M}:=\mathfrak{m}\mathscr{R}(I)+It\mathscr{R}(I)+t^{-1}\mathscr{R}(I)$ which contains $t^{-1}$. By Theorem 3.7, the localization $\mathscr{R}(I)_{\mathfrak{M}}$ is seminormal. Since $G(I)$ is reduced, it follows that the ring $\mathscr{R}(I)$ is also reduced by [18, Exercise 5.9]. Moreover, the integral closure of $\mathscr{R}(I)$ in its total ring of fractions is a $\mathbb{Z}$-graded ring in view of [18, Proposition 5.2.4]. So Theorem 3.6 yields seminormality of $\mathscr{R}(I)$. Since $A$ and $\mathscr{R}_{+}(I)$ are algebra retracts of $\mathscr{R}(I)$ (indeed, $A=[\mathscr{R}(I)]_{0}$ and $\mathscr{R}_{+}(I)=[\mathscr{R}(I)]_{\geq 0}$), they are seminormal as well.

### v2

@@ 1. Introduction
Historically, Rees rings and associated graded rings of Noetherian rings were studied in connection with the problem of resolving singularities of Noetherian schemes by blowing up closed subschemes. The aim of the present article is to study these objects from the ring-theoretic viewpoint. One often encounters a situation where many notable ring-theoretic properties can be derived by studying associated graded rings. Geometrically, associated graded rings correspond to exceptional divisors of a blown-up scheme. Let $A$ be a Noetherian ring with an ideal $I\subseteq A$ and let $\mathscr{R}_{+}(I)$ be its Rees ring. For example, Barshay proved that if $A$ is Cohen-Macaulay and $I$ is generated by a regular sequence, then $\mathscr{R}_{+}(I)$ is also Cohen-Macaulay in [3]. Let $G(I)$ be the associated graded ring of $A$ with respect to $I$. Then Goto and Shimoda characterized the Cohen-Macaulayness/Gorensteinness of $\mathscr{R}_{+}(I)$ in terms of the corresponding property of $G(I)$ and its attached $a$-invariant in [8]. Conversely, one can ask how the singularities of $G(I)$ affect the singularities of $A$. Suppose that $\mathcal{P}$ is one of the following properties: Cohen-Macaulay, Gorenstein, reduced, domain, or normal. Then under a very mild condition, if $G(I)$ has $\mathcal{P}$, then so does $A$ (see [6, Theorem 4.5.7, Theorem 4.5.8 and Theorem 4.5.9] for the proofs). We use Rees rings, extended Rees rings, and associated graded rings to study the relationship among three notable problems in commutative algebra, which are Lifting Problem, Ascent/descent Problem in associated graded rings, and Matijevic-Roberts type Problem. Our guiding principle is explained by Proposition 3.1 and Proposition 3.5. As a test example, we consider the case $\mathcal{P}=\rm{weakly~{}normal,~{}seminormal}$. See Theorem 3.6 and Theorem 3.8.
@@ 2. Notation
As we already mentioned in the introduction, a remarkable aspect of associated graded rings is that one can often derive a certain property of $A$ from $G(I)$. For instance, it is known that if the associated graded ring with respect to some proper ideal of $A$ is Cohen-Macaulay, then $A$ is also Cohen-Macaulay. A similar statement holds for the Gorenstein property (see [13, Proposition 1.2]). Let us recall some definitions. See a survey paper [19] for other equivalent definitions of seminormality and weak normality.
@@
Let $A$ be a commutative reduced Noetherian ring. Consider the following conditions.
@@
A ring $A$ which satisfies the condition $\rm{(i)}$ is called seminormal (see [19, Definition 2.17.]), and which satisfies both conditions $\rm{(i)}$ and $\rm{(ii)}$ is called weakly normal (see [19, Definition 3.12.]).
@@ 1. $\bullet$
    Let $A\in\mathbf{D}$. Then any finitely generated $A$-algebra belongs $\mathbf{D}$, the localization of $A$ with respect to any multiplicative subset of $A$ belongs to $\mathbf{D}$, and any subring $B\subseteq A$ such that $A$ is faithfully flat over $B$ belongs to $\mathbf{D}$. Let $I\subseteq A$ be an ideal. Then the (extended) Rees ring of $A$ with respect to $I$ belongs to $\mathbf{D}$.
We note that if $A\to B$ is faithfully flat and $B$ is Noetherian, then $A$ is also Noetherian. Henceforth, we fix $\mathbf{D}$ as above and consider a ring-theoretic property $\mathcal{P}$ which is defined on any object of $\mathbf{D}$. We consider the following properties on the objects from $\mathbf{D}$.
@@
    If $A$ has $\mathcal{P}$ and $B\to A$ is a faithfully flat extension of rings, then $B$ has $\mathcal{P}$.
@@
by letting $X\mapsto\overline{y}\in yA/y^{2}A$. Then $\phi$ is an isomorphism in view of [11, Theorem 16.2]. Since $A/yA$ has $\mathcal{P}$ by assumption, the condition $(\bf{P2})$ allows us to say that $(A/yA)[X]$ has $\mathcal{P}$. By the graded isomorphism $\phi$, we see that the associated graded ring $G(yA)$ has $\mathcal{P}$. By virtue of assumption $(\bf{Gr})$, it follows that $A$ has $\mathcal{P}$. In other words, $(\bf{Lift})$ is solved. ∎
We recall the following result (see [14, Proposition I.7.4]).
@@
An application of Proposition 3.1 is a new proof of Theorem 3.2, which was already mentioned in [7, Remark 3.8]. However, it seems that most existing research articles study $(\bf{Lift})$ extensively, while $(\bf{Gr})$ is less tractable. As a side topic, we prove an ideal-theoretic analogue of Theorem 3.2 under some conditions.
@@
which is the Rees ring $(A/yA)[(I/(I\cap yA))t]\subseteq(A/yA)[t]$. By assumption, $I/(I\cap yA)\subseteq A/yA$ is a normal ideal and $A/yA$ is integrally closed, so $(A/yA)[(I/(I\cap yA))t]$ is integrally closed in a normal domain $(A/yA)[t]$ in view of [17, Proposition 5.2.1]. Hence $\mathscr{R}_{+}(I)/y\mathscr{R}_{+}(I)$ is an integrally closed domain. Let $\mathfrak{M}:=\mathfrak{m}\oplus\bigoplus_{n\geq 1}I^{n}t^{n}$, which is the unique graded maximal ideal of $\mathscr{R}_{+}(I)$. Since $y\in\mathfrak{M}$, it follows that $\mathscr{R}_{+}(I)_{\mathfrak{M}}$ is a local normal domain. By applying [15, Proposition 2.2], we find that $\mathscr{R}_{+}(I)$ is an integrally closed domain. In particular, this ring is integrally closed in $A[t]$, which gives that $I$ is a normal ideal by [17, Proposition 5.2.1]. ∎
@@
where $\mathscr{R}(I)=A[It,t^{-1}]$ is the extended Rees ring. Note that $\mathscr{R}(I)$ has a unique graded maximal ideal $\mathfrak{M}=\mathfrak{m}\mathscr{R}(I)+It\mathscr{R}(I)+t^{-1}\mathscr{R}(I)$ containing $t^{-1}$. Moreover, $t^{-1}$ is a nonzero divisor of $\mathscr{R}(I)$. By $(\bf{Lift})$ and $(\bf{P1})$, we find that the localization $\mathscr{R}(I)_{\mathfrak{M}}$ has $\mathcal{P}$. Then $(\bf{MR})$ shows that $\mathscr{R}(I)$ has $\mathcal{P}$. By $(\bf{P1})$, the localization $\mathscr{R}(I)[t]=A[t,t^{-1}]$ has $\mathcal{P}$. Since $A[t,t^{-1}]$ is faithfully flat over $A$, the condition $(\bf{P3})$ allows us to conclude that $A$ has $\mathcal{P}$, as desired. ∎
@@
$(1)$: Since seminormality is a local property in view of [16, Proposition 3.7], the “only if” part is clear. It suffices to prove the “if” part. So assume that $A_{\mathfrak{m}}$ is seminormal for every graded maximal ideal $\mathfrak{m}$. Again by [16, Proposition 3.7], it suffices to show that $A_{\mathfrak{p}}$ is seminormal for every maximal ideal $\mathfrak{p}\subseteq A$. Let $\mathfrak{p}^{*}$ be the ideal generated by all homogeneous elements contained in $\mathfrak{p}$. Then it follows from [6, Lemma 1.5.6] that $\mathfrak{p}^{*}$ is a prime ideal. Let $A_{(\mathfrak{p})}$ and $A_{(\mathfrak{p}^{*})}$ be the homogeneous localizations as in [6, p.31]. Then by the construction, we have $A_{(\mathfrak{p}^{*})}=A_{(\mathfrak{p})}$. Since $\mathfrak{p}^{*}$ is graded, there is a graded maximal ideal $\mathfrak{m}$ such that $\mathfrak{p}^{*}\subseteq\mathfrak{m}$. Then we have a localization map $A_{(\mathfrak{m})}\to A_{(\mathfrak{p}^{*})}=A_{(\mathfrak{p})}$, which extends to
@@
However, as the only graded maximal ideal of $A$ is $\mathfrak{m}$ and $b$ is a homogeneous element not contained in $\mathfrak{m}$, we find that $b$ is a unit element of $A$. That is, we get $x\in A$. By applying [10, Proposition 2.4] (see also [1, Theorem 2]), $A$ is seminormal, as desired.
$(2)$: The weak normality is a local property by [20, Corollary 2]. So as in the first case, we may assume that $A$ is a $\mathbb{Z}$-graded ring with a unique maximal ideal $\mathfrak{m}$ and $A_{\mathfrak{m}}$ is weakly normal. Keep the notation as in $(1)$. In view of [18, Proposition 3.1], it suffices to show that if a homogeneous element $x\in\overline{A}$ satisfies $x^{p}\in A$ and $px\in A$ for some prime $p$, then $x\in A$. By the weak normality of $A_{\mathfrak{m}}$, it follows that $\frac{x}{1}\in A_{\mathfrak{m}}$ and one can proceed as in the first case to finish the proof. ∎
@@
While the seminormal case due to Heitmann is found in [9], the weakly normal case due to Murayama is found in [12, Proposition 4.10], where a new proof of the seminormal case is also given. Recently, even another innovative approach is discovered in [4]. The following result gives a realization of Proposition 3.5. The corresponding statement in the normal case has been already known. See for instance [6, Theorem 4.5.9].
@@
$(1)$: First, we prove that the extended Rees ring $\mathscr{R}(I)$ is seminormal. Recall that $G(I)\cong\mathscr{R}(I)/t^{-1}\mathscr{R}(I)$ and $t^{-1}$ is a nonzero divisor of $\mathscr{R}(I)$. Since $A$ is local, $\mathscr{R}(I)$ has a unique graded maximal ideal $\mathfrak{M}:=\mathfrak{m}\mathscr{R}(I)+It\mathscr{R}(I)+t^{-1}\mathscr{R}(I)$ which contains $t^{-1}$. By Theorem 3.7, the localization $\mathscr{R}(I)_{\mathfrak{M}}$ is seminormal. Since $G(I)$ is reduced, it follows that the ring $\mathscr{R}(I)$ is also reduced by [17, Exercise 5.9 at page 116]. Moreover, the integral closure of $\mathscr{R}(I)$ in its total ring of fractions is a $\mathbb{Z}$-graded ring in view of [17, Proposition 5.2.4]. So Theorem 3.6 yields seminormality of $\mathscr{R}(I)$. Then the localization $\mathscr{R}(I)[t]=A[t,t^{-1}]$ is seminormal. Since $A\to A[t,t^{-1}]$ is faithfully flat, it follows that $A$ is seminormal by [19, Theorem 2.22]. The reducedness of $G(I)$ gives that $I$ is a normal ideal, that is $I^{n}$ is integrally closed for all $n>0$ in view of [17, Exercise 5.7 at page 116]. So $\mathscr{R}_{+}(I)$ is integrally closed in the seminormal ring $A[t]$ by [17, Proposition 5.2.1]. Hence $\mathscr{R}_{+}(I)$ is seminormal as well.
@@
Acknowledgement .
Both authors would like to thank the anonymous referee for pointing out errors in the first draft.